OpenStudy (anonymous):

Simplify. –8 + (–6)x – 5 + 2x A. –3 – 4x B. –13 – 4x C. 3 – 8x D. 13 – 8x

jhonyy9 (jhonyy9):

-8-6x-5+2x = (-8-5)+(2x-6x)= -13 -4x B. is right sure

OpenStudy (anonymous):

thanks